# supplement

### v1

@@ 2.4 Quantifying the Impact of Influencers on Secondary Spread
While CRP is a useful indicator of information propagation efficiency, it does not fully capture the scale of that propagation, which can be measured by the actual quantity of reposts. To address this limitation and provide a more comprehensive analysis of influencers’ role in shaping information flow through secondary spread, we investigated the proportion of views and reposts in each influencer category. In our analysis, we calculate the number of times users potentially see shared posts in their simulated timelines, which we refer to as views; these views represent instances where a user would encounter a repost in their timeline (Figure 2(a)). When analyzing the secondary spread in the virtual timeline, it is crucial to distinguish between these views and reposts, as views only indicate exposure to content, whereas reposts represent active propagation of the content to other users. By tracking views, we can determine how often reposts (secondary spread) by users with different influence levels are potentially seen in other users’ timelines. This allows us to quantify not only how many times content is reposted, but also its potential visibility across the repost cascade. Therefore, our approach provides insights into both the spread efficiency (measured by CRP) and the potential reach of information shared by different user categories. Figure 4 illustrates the proportion of users according to influence levels as well as the distribution of views and reposts in secondary spread according to influence categories.
@@
Figure 4: Distribution of secondary spread across influence categories. (a) Proportion of users according to influence levels (for simplicity, percentages are rounded to integers). (b) User view share (proportion of views for each influence category). and user repost share (proportion of reposts that can be traced back to each influence category). The top row shows data for all posts, while the bottom row focuses on posts with $\geq$ 5000 reposts. Notably, very high influence users (the top 1% of the user population) consistently account for approximately half of views and reposts in the secondary spread, demonstrating their substantial impact on information distribution.
@@
The results reveal that users with very high influence, despite comprising only 1% of the user base (Figure 4(a)), account for 58.0% of views and 53.3% of reposts in secondary spread across all posts (Figure 4(b) above). This disproportionate influence becomes somewhat less pronounced when the number of reposts is large ($\geq$ 5000), where very high influence users are responsible for 40.6% of views and 47.7% of reposts (Figure 4(b) bottom). Notably, for highly popular posts ($\geq$ 5000 reposts), an interesting shift occurs in the behavior of very high influence users. While their overall share of both views and reposts decreases compared to all posts, their share of reposts (47.7%) now exceeds their share of views (40.6%). This contrasts with the pattern observed for all posts, where their share of views (58.0%) is higher than their share of reposts (53.3%). This reversal is particularly significant, given that reposts can only occur after a user views the content. This finding not only aligns with the high CRP observed for this category in Figure 3(b) but also indirectly demonstrates that when influencers share content, it generally receives high engagement from their followers, thereby contributing to its popularity. In this way, the role of very high influence users shifts with regards to viral content. Instead of merely exposing content to their large follower base, they are effectively amplifying the content through their own sharing (by encouraging reposting)). This amplifying effect of influencers plays a crucial role in accelerating and expanding the propagation of popular content through secondary spread.
2.5 Summary and Interpretation of Influencers’ Impact on Information Diffusion
Our analysis combining CRP with the quantification of view and repost shares provides a comprehensive picture of how user influence impacts information diffusion in online social networks. These results support our hypothesis regarding prestige bias and reveal the extent to which a small group of highly influential users shape information flow. As hypothesized, our findings align with the concept of prestige bias in several key ways:
  1. 1.
    Users with very high influence consistently demonstrate higher CRP over time for popular posts (Figure 3) for popular posts.
  2. 2.
    Very high influence users (1% of user base) account for 53.3% and 47.7%of reposts in secondary spread across all posts and posts with $\geq$ 5000 repost, respectively (Figure 4(b) above).
  3. 3.
    For highly popular posts ($\geq$ 5000 reposts), very high influence users’ share of reposts (47.7%) exceeds their share of views (40.6%) (Figure 4(b) bottom).
  4. 4.
    Very high influence users maintain substantially higher CRP throughout the 24 hour observation period (Figure 3(b)).
Importantly, these effects are not due to more frequent reposting by highly influential users. Analysis in Supplementary Material LABEL:sup-reposting-behavior demonstrates that very high influence users do not repost more often than users in other influence categories. Instead, the impact comes from the wider reach of influential users’ reposts and the higher likelihood of those reposts being shared further (higher CRP). This confirms that prestige bias manifests through enhanced diffusion of content shared by influencers. As discussed above, for highly popular posts ($\geq$ 5000 reposts), we observe a notable shift in very high influence users’ impact; their share of reposts exceeds their share of views, reversing the pattern seen in all posts (Figure 4 (b)). This reversal in the view-to-repost ratio for viral content suggests that influential users become even more effective at driving engagement as content gains popularity, indicating that prestige bias plays a particularly potent role in amplifying the spread of viral information. To further investigate this phenomenon, we conducted additional analyses on the relationship between user influence and virality, which can be found in Supplementary Material LABEL:sup-viral-diffusion. In summary, our results provide strong evidence for the effect of prestige bias in online information sharing. Content reposted by influential users consistently reaches a broader audience and is diffused more widely, supporting our initial hypothesis. These findings highlight the significant role that a small group of highly influential users play in shaping information diffusion patterns on social media platforms. This has important implications for understanding and potentially managing information diffusion in digital environments.
@@ 3 Discussion
In this study, we introduced new concepts (secondary spread and CRP) to test the hypothesis that influencers (users with high hg-index scores) more effectively propagate information when resharing others’ content than non-influencers; this analysis was conducted using one month of data Japanese-language posts on Twitter (currently X).
@@
Our analysis revealed that influencers consistently demonstrate significantly higher CRP for popular posts ($\geq$ 1,000 reposts) in secondary spread compared to non-influencers. The CRP indicating influencers’ impact remained high over time, showing a robust sustained influence on information diffusion. Moreover, influencers were found to substantially impact the actual distribution of reposts in secondary spread. Over half of the views and reposts in secondary spread could be attributed to the top 1% of users, with this trend being even more pronounced for highly popular posts. One might argue that the CRP value is inherently low in absolute terms (e.g., 0.01). However, this is likely because we count all reposts appearing in the virtual timeline as views. If we could evaluate only the posts that users actually viewed, the probability might be slightly higher. Nevertheless, as suggested by previous studies  [18, 20], consecutive reposts occurring in sequence remains a rare phenomenon. Even if the absolute term of the CRP is low, its relatively high value in the early stages can still be considered a crucial metric if it influences the subsequent size of the cascade for that post.
@@
Our findings provide strong evidence for the effect of prestige bias in online social networks. Users with higher influence, as measured by their hg-index, consistently demonstrate a greater ability to propagate information, even when that information is not their own original content. This effect persists throughout the important early stages of information diffusion. These results align with the theoretical framework of prestige bias proposed in anthropology and psychology [13, 14], suggesting that this cognitive tendency extends to digital environments. Influential users’ consistently higher CRP values indicate that their status enhances their ability to spread information, supporting our initial hypothesis. The novelty of this research lies in its integration of the roles of influencers as source spreaders (originators of content) and brokers (information intermediaries), which have been the focus of previous influencer studies. By introducing the concepts of secondary spread and CRP, we have revealed how influencers effectively fulfill both roles, providing a new perspective that bridges existing research areas. Furthermore, this study empirically demonstrates the effect of prestige bias in online communities. We have shown that information reposted by influencers tends to spread more widely and persistently.
@@
Interestingly, influencers did not demonstrate much influence in relation to unpopular posts. This finding suggests that prestige bias may depend both on the influencer and content characteristics. Traditional theories assumed that prestige bias uniformly affected all kinds of information; our results suggest this may not be the case in online communities. In this regard, the research by Acerbi and Tehrani [24] provides important insights. Their experiments showed that in selecting quotations the content was more important than their attribution to famous individuals. This result suggests that in online environments, the quality of content may be more pertinent in evaluating information than the status of the sender. However, our results also suggest that he prestige of the sharer information may becomes important when content has strong appeal. That is, prestige bias may have a stronger role when the quality of content is equivalent. Moreover, while previous research such as that by Brand et al. [25] has shown that source expertise influences the perceived reliability of information on social media, our findings significantly extend this understanding.
@@
As detailed in Supplementary Material LABEL:sup-viral-diffusion, our analysis of structural virality [26] in widely diffused posts reveals that influencers’ impact actively shapes information diffusion patterns. Posts reposted by influencers not only reach a wider audience but also exhibit higher structural virality, indicating more complex and extensive diffusion. This, coupled with the high CRP observed among influencers for potentially viral content, suggests a dual role for influencers: they act as both broadcasters, directly reaching large audiences, and amplifiers, significantly increasing the likelihood of further sharing by their followers. By enhancing the shareability of their posts (and reposts), influencers create a ripple effect extending beyond their direct connections. Identifying this amplifying role in viral diffusion represents a significant advance in our understanding of social media information spread, moving beyond simple wide-reach effects to a more nuanced view of influencers’ impact on content propagation.
@@
This research provides a bridge linking cultural evolution to a series of studies on online information diffusion from a computational social science perspective. Cultural evolution research has presented many insights into the adaptive aspects of human cognition through psychological experiments and mathematical simulations. Meanwhile, information diffusion research has presented quantitative analyses of the mechanisms of misinformation spread and echo chamber phenomena through analysis of large-scale data from social network services. By integrating these two research lines, our study confirmed that adaptive theories relating to and hypotheses of cognition presented in cultural evolution theory can be observed even in social networks, a contemporary digital environment for human relationships. Furthermore, we demonstrated that prestige bias manifests in online communities, where information can be transmitted to others through easy sharing functionalities.
Our results emphasize the importance of influencers in marketing on social media [17, 2, 18, 5]. The fact that influencers’ influence is stronger for certain types of content (i.e., influencers’ impact on secondary spread varies depending on the content) can aid decision-making around influencer marketing strategies. These results implicitly suggest that influencers play a role in the spread of misinformation, which has become a significant social issue. In recent years, research has attempted to propose solutions to the issue of misinformation from the perspectives of education and technology [27] and behavioral science [28]. Notably, it has been suggested that influencers (especially experts) can be effective in correcting misinformation [7]. The present study complements this finding (which addresses the perspective of the sender or influencer) from the perspective of the receiver; our results suggest that information resharing (secondary spread) by influencers can also play an important role in the correction of misinformation. Notably, the study by Lim et al. [7] dealt with a limited case of correcting specific medical terms, and caution is needed in generalizing this to the correction of broader misinformation or information diffusion in general. Nevertheless, our results suggest that it is important to effectively utilize the power of influencers when addressing misinformation, and provide useful hints for narrowing down the targets of the application. This approach may enhance the efficiency and effectiveness of misinformation countermeasures.
@@
In conclusion, this study empirically demonstrated the existence and function of prestige bias in online communities and revealed that influencers have a significant impact on information diffusion through secondary spread. These findings make important contributions to understanding the dynamics of information diffusion in social media and the role of influencers. The concepts of primary and secondary spread proposed in this study, as well as the framework using virtual timeline and repost cascades, provide new perspectives for research on information diffusion in online communities. This research deepens our understanding of the influence of social media and provides insights into the complex interactions between user status, content sharing, and information propagation in online spaces. These findings can inform various applications, such as the development of more effective information diffusion strategies and measures against misinformation spread.
@@ 4.1 Data Collection and Preprocessing
We sampled Japanese-language reposts from Twitter (currently X) using the streaming API from October 1 to October 1 to 31, 2021. We chose Japanese posts as the object of analysis for the following reasons:
@@
    A more limited and homogeneous sample compared to English-language posts
Our one-month data collection period aligns with timeframes used in previous research [18, 9]. We chose this timeframe for several reasons. First, our large-scale sampling method enabled us to gather a volume of data comparable to that of studies that utilized several years’ worth of information. Secondly, considering that the primary focus of this research is on the relationship between influencers and instantaneous virality, a one-month period provides a sufficiently comprehensive snapshot to capture these dynamics.
To reconstruct follower-followee relationships as of October 2021, we sampled users every other day during September 2021, retrieved their following and follower lists, and synthesized these snapshots. Notably, we were unable to capture private accounts, for which follower/following information is not publicly accessible. Our network reconstruction is thus limited to publicly visible connections and provides an approximation of the actual user links during the period; there may be some discrepancies with regards to actual follower-followee relationships.
Notably, our analysis focused solely on simple reposts in Twitter and excluded quote posts. This limitation is an important point to consider when interpreting the results.

### v3

@@ 2.5 Quantifying the Impact of Influencers on Secondary Spread
Although the CRP is a useful indicator of information propagation efficiency, it does not fully capture the scale of that propagation, which can be measured by the actual quantity of reposts. To address this limitation and provide a more comprehensive analysis of influencers’ role in shaping information flow through secondary spread, we investigated the proportion of views and reposts in each influencer category. In our analysis, we calculate the number of times users potentially see shared posts in their simulated timelines, which we refer to as views; these views represent instances where a user would encounter a repost in their timeline (Figure 2(a)). When analyzing the secondary spread in the virtual timeline, it is crucial to distinguish between these views and reposts, as views only indicate exposure to content, whereas reposts represent active propagation of the content to other users. By tracking views, we can determine how often reposts (secondary spread) by users with different influence levels are potentially seen in other users’ timelines. This allows us to quantify not only how many times content is reposted but also its potential visibility across the repost cascade. Therefore, our approach provides insights into both the spread efficiency (measured by the CRP) and the potential reach of information shared by different user categories. Figure 5 illustrates the proportion of users according to influence levels as well as the distribution of views and reposts in the secondary spread according to influence categories.
@@
Figure 5: Distribution of secondary spread across influence categories. (a) Proportion of users according to influence levels (for simplicity, percentages are rounded to integers). (b) User view share (proportion of views for each influence category). and user repost share (proportion of reposts that can be traced back to each influence category). The top row shows data for all posts, whereas the bottom row focuses on posts with $\geq$ 5000 reposts. Notably, very high-influence users (the top 1% of the user population) consistently account for approximately half of the views and repost in the secondary spread, demonstrating their substantial impact on information distribution.
@@
The results reveal that users with very high influence, despite comprising only 1% of the user base (Figure 5(a)), account for 58.0% of views and 53.3% of reposts in secondary spread across all posts (Figure 5(b) above). This disproportionate influence becomes somewhat less pronounced when the number of reposts is large ($\geq$ 5000), where very high-influence users are responsible for 40.6% of views and 47.7% of reposts (Figure 5(b) bottom). Notably, for highly popular posts ($\geq$ 5000 reposts), an interesting shift occurs in the behavior of very high-influence users. Although their overall share of both views and reposts decreases compared with all posts, their share of reposts (47.7%) now exceeds their share of views (40.6%). This contrasts with the pattern observed for all posts, where their share of views (58.0%) is greater than their share of reposts (53.3%). This reversal is particularly significant, given that reposts can only occur after a user views the content. This finding not only aligns with the high CRP observed for this category in Figure 3(b) but also indirectly demonstrates that when influencers share content, it generally receives high levels of engagement from their followers, thereby contributing to its popularity. In this way, the role of very high-influence users shifts with respect to viral content. Instead of merely exposing content to their large follower base, they effectively amplify the content through their own sharing (by encouraging reposting). This amplifying effect of influencers plays a crucial role in accelerating and expanding the propagation of popular content through secondary spread.
2.6 Influence of the First-Reposted Users on Viral Diffusion
Thus far, we have focused only on the direct followers of influencers; however, influencers can further affect nondirect followers (followers’ followers) through a repost cascade. Here, we further analyze the depth and complexity of the repost cascade.
An analysis of repost cascades for posts that attracted more than 5,000 reposts revealed a strong relationship between the influence level of the first user who shared a post and its diffusion depth and complexity. Figure 6 illustrates the conceptual structure of a repost subtree, with each rectangular layer representing a stage in the diffusion process. When high- or very high-influence users make the initial repost, the content typically propagates through more tiers of sharing, resulting in greater structural virality and an increased maximum depth.
Structural virality was measured by calculating the average path length among all pairs of users within each repost cascade. This analysis provided a sense of how extensively and how many steps the content traveled. The maximum depth was determined by measuring the longest path from the first-reposted users to the most distant individual in the repost chain. As shown in Figure 7, both metrics tended to increase with increasing user influence. In particular, the very high category presented the largest values, indicating that these individuals—often characterized as influencers—did more than simply broadcast posts to a large audience. They appeared to promote multistep diffusion cascades, suggesting that followers who received the post from a highly influential source were more inclined to share it further, extending its overall reach and complexity. These findings underscore the importance of accounting for user influence in viral diffusion.
Figure 6: Conceptual repost subtree rooted at the first-reposted users. Rectangular regions represent progressively deeper layers of sharing.
Figure 7: (Left) structural virality and (right) maximum depth across different user influence levels. The error bars represent confidence intervals for each influence category.
@@ 3 Discussion
In this study, we introduced new concepts (secondary spread and CRP) to test the hypothesis that influencers (users with high hg index scores) more effectively propagate information when resharing others’ content than noninfluencers do; this analysis was conducted using one month of Japanese-language post data.
@@
Our analysis revealed that influencers consistently demonstrate significantly higher CRP for popular posts ($\geq$ 1,000 reposts) in secondary spreads than noninfluencers. The CRP, indicating influencers’ impact, remained high over time, showing a robust, sustained influence on information diffusion. Moreover, influencers were found to substantially impact the actual distribution of reposts in secondary spread. Over half of the views and reposts in secondary spreads could be attributed to the top 1% of users, with this trend being even more pronounced for highly popular posts. One might argue that the CRP value is inherently low in absolute terms (e.g., 0.01). However, this is likely because we count all reposts appearing in the virtual timeline as views. If we could evaluate only the posts that users actually viewed, the probability might be slightly higher. Nevertheless, as suggested by previous studies  [18, 20], consecutive reposts occurring in sequence remain rare phenomena. Even if the absolute term of the CRP is low, its relatively high value in the early stages can still be considered a crucial metric if it influences the subsequent size of the cascade for that post. Note that these effects are not due to more frequent reposting by highly influential users. The analysis in Supplementary Material S4 demonstrates that very high-influence users do not repost more often than users in other influence categories. Instead, the impact comes from the wider reach of influential users’ reposts and the greater likelihood of those reposts being shared further (higher CRP). Moreover, these results were replicated on the one-day post dataset of the English language.
A possible concern is that relying solely on a strictly chronological sequence may overlook the impact of engagement and advertising. We restricted our analysis to the first 30 min after a post’s publication to reduce the influence of algorithmic recommendations. Many early-stage posts are unlikely to have significant differences in engagement, which may help mitigate the influence of algorithms. The results (Supplementary Material S5) remained consistent with those reported here, reinforcing the reliability of our virtual timeline approach.
Additionally, the results of the mixed effects logistic regression controlling for post-level and user-level heterogeneity were consistent with the results of the CRPs in subsection 2.2.
@@
Our findings provide consistent evidence for the effect of prestige bias in online social networks. Users with greater influence, as measured by their hg index, consistently demonstrate a greater ability to propagate information, even when that information is not their own original content. This effect persists throughout the important early stages of information diffusion. These results align with the theoretical framework of prestige bias proposed in anthropology and psychology [13, 14], suggesting that this cognitive tendency extends to digital environments. Influential users’ consistently higher CRP values indicate that their status enhances their ability to spread information, supporting our initial hypothesis. The novelty of this research lies in its integration of the roles of influencers as source spreaders (originators of content) and brokers (information intermediaries), which have been the focus of previous influencer studies. By introducing the concepts of secondary spread and the CRP, we reveal how influencers effectively fulfill both roles, providing a new perspective that bridges existing research areas. Furthermore, this study empirically demonstrates the effect of prestige bias in online communities. We have shown that information reposted by influencers tends to spread more widely and persistently.
@@
Interestingly, influencers did not demonstrate much influence in relation to unpopular posts. This finding suggests that prestige bias may depend on both influencer and content characteristics. Traditional theories assume that prestige bias uniformly affects all kinds of information; our results suggest that this may not be the case in online communities. This implication is also consistent with the results that the variance of source_tweet_id is greater than the user_influence coefficients of very high in the mixed-effects logistic regression analysis, which is detailed in Supplementary Material S3. In this context, the research by Acerbi and Tehrani [26] provides important insights. Their experiments revealed that in selecting quotations, the content was more important than their attribution to famous individuals. This result suggests that in online environments, the quality of content may be more pertinent in evaluating information than the status of the sender. However, our results also suggest that the content of the shared information may become important when the content has a strong appeal. That is, prestige bias may play a stronger role when the quality of content is equivalent. Moreover, although previous research, such as that by Brand et al. [27], has shown that source expertise influences the perceived reliability of information on social media, our findings significantly extend this understanding.
@@
Our analysis of structural virality [28] in widely diffused posts reveals that influencers’ impact actively shapes information diffusion patterns. Posts reposted by influencers not only reach a wider audience but also exhibit greater structural virality, indicating more complex and extensive diffusion. This, coupled with the high CRP observed among influencers for potentially viral content, suggests a dual role for influencers: they act as both broadcasters, directly reaching large audiences, and amplifiers, significantly increasing the likelihood of further sharing by their followers. By enhancing the shareability of their posts (and reposts), influencers create a ripple effect extending beyond their direct connections. Identifying this amplifying role in viral diffusion represents a significant advance in our understanding of social media information spread, moving beyond simple wide-reach effects to a more nuanced view of influencers’ impact on content propagation. However, one possible explanation for these findings is that we have defined influencers via the hg index. Indeed, the results of the mixed-effects logistic regression (Table 2) suggest that sender_follower_counts is not necessarily correlated with sender_influence. This result indicates that a larger follower count does not automatically equate to greater influence (as operationalized by the hg index). Instead, these two measures seem to capture different facets of user impact, underscoring the importance of distinguishing between raw follower counts and influence metrics when evaluating an individual’s role in information diffusion.
@@
This research provides a bridge linking cultural evolution to a series of studies on online information diffusion from a computational social science perspective. Cultural evolution research has presented many insights into the adaptive aspects of human cognition through psychological experiments and mathematical simulations. Moreover, information diffusion research has presented quantitative analyses of the mechanisms of false information spread and echo chamber phenomena through the analysis of large-scale data from social network services. By integrating these two lines of research, our study confirmed that adaptive theories related to and hypotheses of cognition presented in cultural evolution theory can be observed even in social networks, a contemporary digital environment for human relationships. Furthermore, we demonstrated that prestige bias manifests in online communities, where information can be transmitted to others through easy sharing functionalities.
Our results emphasize the importance of influencers in marketing on social media [17, 2, 18, 5]. The fact that influencers’ influence is stronger for certain types of content (i.e., influencers’ impact on secondary spread varies depending on the content) can aid decision-making around influencer marketing strategies. These results implicitly suggest that influencers play a role in the spread of misinformation, which has become a significant social issue.
Additionally, in recent years, researchers have attempted to propose solutions to the issue of false information from the perspectives of education and technology [29] and behavioral science [30].
There is an argument that the spread of false information online is related to human behavior and cognition [31], and research has also shown that influencers can play an effective role in correcting false information [7]. Given these considerations, the prestige bias highlighted in this study can be linked to one of the cognitive biases that influence information diffusion. From this perspective, this study may have potential applications in research on actual false information.
Nevertheless, as this study demonstrates, the content of a post is also essential for reposting. Although this study did not distinguish between false information and other types of information, future research is needed to investigate how the correctness of information and cognitive biases interact.
However, the content is still important. Therefore, we need to research interactions with content in future work.
@@
In conclusion, this study empirically demonstrated the existence and function of prestige bias in online communities and revealed that influencers have an impact on information diffusion through secondary spread. The concepts of primary and secondary spread proposed in this study, as well as the framework using virtual timelines and repost cascades, provide new perspectives for research on information diffusion in online communities. This research deepens our understanding of the influence of social media and provides insights into the complex interactions among user status, content sharing, and information propagation in online spaces. These findings can inform various applications, such as the development of more effective information diffusion strategies and measures against false information spread.
@@ 4.1 Data Collection and Preprocessing
We sampled Japanese-language reposts from Twitter (currently X) via the streaming API from October 1 to October 1 to 31, 2021. We chose Japanese posts as the objects of analysis for the following reasons:
@@
    A more limited and homogeneous sample than English-language posts
Our one-month data collection period aligns with the timeframes used in previous research [18, 9]. We chose this timeframe for several reasons. First, our large-scale sampling method enabled us to gather a volume of data comparable to that of studies that utilized several years of information. Second, considering that the primary focus of this research is the relationship between influencers and instantaneous virality, a one-month period provides a sufficiently comprehensive snapshot to capture these dynamics.
To reconstruct follower–followee relationships as of October 2021, we used sampled users from users who had reposted Japanese posts, and we retrieved their following and follower lists every other day during September 2021. Notably, we were unable to capture private accounts, for which follower/following information is not publicly accessible. Our network reconstruction is thus limited to publicly visible connections and provides an approximation of the actual user links during the period; there may be some discrepancies with respect to actual follower‒followee relationships.
Notably, our analysis focused solely on simple reposts on Twitter and excluded quote posts. This limitation is an important point to consider when interpreting the results.